Let W be a solutions space of homogeneous equation 2a+3b+4c=0
Describe the cosets of W in R^3
1
Expert's answer
2020-12-22T18:45:20-0500

"Solution: Given ~that~ "W~ be~ a~ solution~ space~ of~ homogenuous~ equation~ 2a+2b+4c=0""

"\therefore solving ~2a+2b+4c=0
\\ \therefore a+b+2c=0
\\a=-b-2c~where ~b ~and~c~ are~ free ~variables.
\\\ This ~can~be ~written~as~"

"W=\begin{bmatrix}
a \\
b \\
c\\
\end{bmatrix}={\begin{bmatrix}
-b-2c \\
b \\
c\\
\end{bmatrix}}=-b\begin{bmatrix}
-1 \\
1\\
0\\
\end{bmatrix}+c\begin{bmatrix}
-2\\
0\\
1\\
\end{bmatrix}"

"\therefore W=\begin{bmatrix}
-1 \\
1\\
0\\
\end{bmatrix},\begin{bmatrix}
-2\\
0\\
1\\
\end{bmatrix}
~ these ~are~ the~ two~vectors~in~R^3"

"now~ after~ observing ~theses~ vectors~we ~conclude~ that~,these ~are~equations~in~straight~lines." In three dimensional space the cosets of a line (plane) through the origin is the set of all lines (planes) parallel to it.

If we consider the equation "a=-b-2c~where ~b ~and~c~ are~ free ~variables."

just replacing b and c values, keeping a constant, we get different solutions and all these solutions are gives parallel lines to each other. So all these are cosets of the given line 2a+2b+4c=0 in "R^3."